1. An employee receives an hourly wage rate of $24, with time-and-a-half for all hours worked in excess of 40 during the week. Payroll data for the current week are as follows: hours worked, 44; federal income tax withheld, $145; social security tax rate, 6.0%; Medicare tax rate, 1.5%; state unemployment compensation tax, 3.4% on the first $7,000; and federal unemployment compensation tax, 0.8% on the first $7,000. What is the net amount to be paid to the employee? Round your answer to the nearest cent.

2. Thomas Martin receives an hourly wage rate of $19, with time-and-a-half pay for all hours worked in excess of 40 hours during a week. Payroll data for the current week are as follows: hours worked, 46; federal income tax withheld, $347; social security tax rate, 6.0%; and Medicare tax rate, 1.5%. What is the gross pay for Martin?
Answer #1

Part 1

Regular pay (40*24)

960

Overtime pay ((44-40)*24*1.5)

144

Gross pay

$1104

federal income tax withheld = 145

Social security tax = 1104*6% = 66.24

Medicare tax = 1104*1.5%

Net pay = gross pay - federal income tax withheld - Social security tax - Medicare tax = 1104-145-66.24-16.56 = $876.20

Part 2

Regular pay (40*19)

760

Overtime pay ((46-40)*19*1.5)

171

Gross pay

$931